There is a NACA airfoil with a mean camber line that is below:

z/c = 0.600 [0.5(x/c)-(x/c)^2] for 0\leq x/c\leq 0.25

z/c = 0.111[0.3 +0.4(x/c)-(x/c)^2] for 0.25\leq x/c\leq 1.0

Given thin airfoil theory, find

a) the angle of attack at zero lift

b) the lift coefficient when \alpha =5 degrees
Answer #1
  1. The above problem is related to airfoil theory. Lift coefficient and angle of attack is evaluated.
  2. All the necessary steps are clearly mentioned and easy to understand.
  3. Hope the answer will help you. Thank you for your support.
  4. Given T NACA is below: z Now Using the (a) The Solution i airfoil with camber line (o that 0.600(0.5(*)-()*] for 0x75 0.25 0.
  5. Now, Putting xa (1 - Coso) 0.6 (1-coso) for oLoL 1.047 0.3 dz da 0.0444 - 0.111 (1-(050) for 1.047 4OLIT lah -0.3 + 0,6 Coso
  6. (6) The lift coefficient when a=5 degrees C, = 21 (K-2) Cu = 21 ( 5+ 6.64) x- 1.C, = 1.276 ग 180°